OpenStudy (anonymous):

f(x)= 2x-4 find -f(x)?

jimthompson5910 (jim_thompson5910):

Similar example: f(x)= 5x-7 -f(x)= -( 5x-7) -f(x)= -5x+7 So the answer is -f(x)= -5x+7

OpenStudy (anonymous):

so my answer is -2x + 4?

jimthompson5910 (jim_thompson5910):

it is

OpenStudy (anonymous):

thanks for the example

jimthompson5910 (jim_thompson5910):

np